# Explaining Similarity in Vision–Language Encoders with Weighted Banzhaf Interactions

### v1

@@ C.1 Models
  • CLIP ViT-B/32 [Radford et al., 2021]: openai/clip-vit-base-patch32 (MIT License)
  • CLIP ViT-B/16 [Radford et al., 2021]: openai/clip-vit-base-patch16 (MIT License)
  • SigLIP ViT-B/16 [Zhai et al., 2023]: google/siglip-base-patch16-224 (Apache License 2.0)
  • SigLIP ViT-L/16 [Zhai et al., 2023]: google/siglip-large-patch16-256 (Apache License 2.0)
  • SigLIP-2 ViT-B/32 [Tschannen et al., 2025]: google/siglip2-base-patch32-256 (Apache License 2.0)
  • SigLIP-2 ViT-B/16 [Tschannen et al., 2025]: google/siglip2-base-patch16-224 (Apache License 2.0)
  • SigLIP-2 ViT-L/16 [Tschannen et al., 2025]: google/siglip2-large-patch16-256 (Apache License 2.0)
We additionally rely on the CLIP ViT-B/32 and ViT-B/16 model versions from the clip Python library [Radford et al., 2021, MIT License], not to modify the official implementation of the baselines (Appendix C.3). We set the batch size to 64 for the base models and to 32 for the large models, performing computation on A100 GPUs with 40GB VRAM.
@@ C.2 Datasets
  • MS COCO test set [Lin et al., 2014]: clip-benchmark/wds_mscoco_captions (CC BY 4.0)
  • ImageNet-1k validation set [Deng et al., 2009]: ILSVRC/imagenet-1k (ImageNet Agreement)
@@
Regarding ImageNet-1k, we use all 50 images from each of the following 10 class labels for constructing the pointing game evaluation: goldfish (1), husky (248), cat (282), plane (404), church (497), ipod (605), ball (805), tractor (866), banana (954), pizza (963). We combine these images with labels into 10 varied games as follows: goldfish-husky-pizza-tractor, cat-goldfish-plane-pizza, banana-cat-tractor-ball, husky-banana-plane-church, pizza-ipod-goldfish-banana, ipod-cat-husky-plane, tractor-ball-banana-ipod, plane-church-ball-goldfish, church-pizza-ipod-cat, ball-husky-banana-tractor. In each case, we distinguish four scenarios by gradually adding each consecutive token from left to right, which leads to 40 scenarios of 50 images total, for a reasonable case.
@@ C.3 Baselines
We were not able to run XCLIP on the CLIP (ViT-B/16) model version for inputs with more than 24 text tokens, which are common in MS COCO, running into out-of-memory errors. Furthermore, we refrain from attempting to run these implementations on SigLIP models, which were not part of the original experiments. For context, we did not find an openly available implementation of InteractionLIME [Joukovsky et al., 2023], while it is not clear how to adapt InteractionCAM [Sammani et al., 2024] (proposed for image–image encoders) to the text/language domain.
@@ C.5 Compute resources
Experiments described in Section 5 and Appendix D were computed on a cluster consisting of $4\times$ AMD Rome 7742 CPUs (256 cores), 4TB of RAM, and $16\times$ A100 GPUs for about 16 days combined. We envision that preliminary and failed experiments required the same amount of compute resources.
@@ Appendix D Additional Experimental Results
Figure 10: Insertion/deletion curves for CLIP (ViT-B/16) on MS COCO. Extended Figure 3. AID score (higher is better) for FIxLIP against alternative explanation methods, where a random baseline scores 0. The y-axis is normalized between the model’s prediction on the original input ($100\%$) and the fully removed one ($0\%$), where negative values denote that the model is predicting the inputs are unsimilar. Gradient-based methods like Grad-ECLIP fail to recover nonlinear rankings of important tokens, while our method faithfully recovers the optimal subset explanation. XCLIP does not appear in the comparison as its official implementation returns an OOM error when querying the larger model ($14\times 14$ image tokens) with texts longer than 24 tokens, which are common in MS COCO.
@@
Table 4: Pointing game recognition for CLIP (ViT-B/16) on ImageNet-1k. Extended Table 1. PGR score for FIxLIP against alternative explanation methods, where a random baseline scores $0.25$. The 1st-order methods like Grad-ECLIP and Shapley values fail to distinguish between multiple objects at once, while 2nd-order methods faithfully recover the appropriate explanation (up to the pointing game’s irreducible non-optimality).
@@
GAME [Chefer et al., 2021] $60_{\pm.11} .$$41_{\pm.04} .$$33_{\pm.02} .$$27_{\pm.01}$
Grad-ECLIP [Zhao et al., 2024] $72_{\pm.13} .$$44_{\pm.04} .$$33_{\pm.01} .$$26_{\pm.01}$
@@ E.1 Guidance on interpreting visualizations of interaction explanations
Finding interesting subsets. FIxLIP explanations allow searching for interesting subsets $M$ like $M_{\mathbf{e},\max,k}$ or $M_{\mathbf{e},\min,k}$ for varying sizes of $k$. We can search the space of interactions and visualize only the explanations for tokens in these subsets, or overlay the masks on top of the input. In Figure 13D, we illustrate $M_{\mathbf{e},\max,7}$ containing tokens related to elephant. Starting the search from the bird token retrieves the $k=4$ subset with information about the bird. While the elephant subset increases the similarity, the bird subset could slightly decrease the similarity, presumably because the model would find the isolated image–text similarity of the elephant more clear.
@@ E.2 Additional examples
Figure 14 demonstrates the utility of conditioning on tokens. Figure 15 shows intriguing examples of interactions in inputs with text written in an image.
@@ E.3 Comparison between CLIP (ViT-B/32) and SigLIP-2 (ViT-B/32)
See Figure 16. Note that these models have different text tokenizers and input image resolutions, i.e. $224\times 224$ vs. $256\times 256$, which is why the inputs differ slightly.
@@ E.4 Comparison between CLIP (ViT-B/32) and CLIP (ViT-B/16)
Visual examples between CLIP (ViT-B/32) and CLIP (ViT-B/16) are in Figure 17. Notably, in the example including the clock (Figure 17, row 3), conditioning on the vision token containing the letters T and part of O reveals a strong interaction with another vision token of this word in the “sign” and the text token town. This may suggest the presence of a third-order interaction between these tokens.

### v2

@@ C.1 Models
  • CLIP ViT-B/32 [radford2021learning]: openai/clip-vit-base-patch32 (MIT License)
  • CLIP ViT-B/16 [radford2021learning]: openai/clip-vit-base-patch16 (MIT License)
  • SigLIP ViT-B/16 [zhai2023siglip]: google/siglip-base-patch16-224 (Apache License 2.0)
  • SigLIP ViT-L/16 [zhai2023siglip]: google/siglip-large-patch16-256 (Apache License 2.0)
  • SigLIP-2 ViT-B/32 [tschannen2025siglip2]: google/siglip2-base-patch32-256 (Apache License 2.0)
  • SigLIP-2 ViT-B/16 [tschannen2025siglip2]: google/siglip2-base-patch16-224 (Apache License 2.0)
  • SigLIP-2 ViT-L/16 [tschannen2025siglip2]: google/siglip2-large-patch16-256 (Apache License 2.0)
We additionally rely on the CLIP ViT-B/32 and ViT-B/16 model versions from the clip Python library [radford2021learning, MIT License], not to modify the official implementation of the baselines (Appendix C.3). We set the batch size to 64 for the base models and to 32 for the large models, performing computation on A100 GPUs with 40GB VRAM.
@@ C.2 Datasets
  • MS COCO test set [lin2014microsoft]: clip-benchmark/wds_mscoco_captions (CC BY 4.0)
  • ImageNet-1k validation set [deng2009imagenet]: ILSVRC/imagenet-1k (ImageNet Agreement)
@@
Regarding ImageNet-1k, we use all 50 images from each of the following 10 class labels for constructing the pointing game evaluation: goldfish (1), husky (248), cat (282), plane (404), church (497), ipod (605), ball (805), tractor (866), banana (954), pizza (963). We combine these images with labels into 10 varied games as follows: goldfish-husky-pizza-tractor, cat-goldfish-plane-pizza, banana-cat-tractor-ball, husky-banana-plane-church, pizza-ipod-goldfish-banana, ipod-cat-husky-plane, tractor-ball-banana-ipod, plane-church-ball-goldfish, church-pizza-ipod-cat, ball-husky-banana-tractor. In each case, we distinguish four scenarios by gradually adding each consecutive token from left to right, resulting in 40 scenarios, each with a total of 50 images, which constitutes a reasonable case.
@@ C.3 Baselines
We were unable to run exCLIP on the CLIP (ViT-B/16) model version for inputs with more than 24 text tokens, which are common in MS COCO, running into out-of-memory errors. Furthermore, we refrain from attempting to run these implementations on SigLIP models, which were not part of the original experiments. For context, we did not find an openly available implementation of InteractionLIME [joukovsky2023model], while it is not clear how to adapt InteractionCAM [sammani2024visualizing] (proposed for image–image encoders) to the text/language domain.
@@ C.5 Compute resources
Experiments described in Section 5 and Appendix D were computed on a cluster consisting of $4\times$ AMD Rome 7742 CPUs (256 cores), 4TB of RAM, and $16\times$ A100 GPUs for about 15 days combined. We envision that preliminary and failed experiments have required the same amount of compute resources.
@@ Appendix D Additional Experimental Results
Figure 10: Insertion/deletion curves for CLIP (ViT-B/16) on MS COCO. Extended Figure 3. AID score (higher is better) for FIxLIP against alternative explanation methods, where a random baseline scores 0. The y-axis is normalized between the model’s prediction on the original input ($100\%$) and the fully removed one ($0\%$), where negative values denote that the model is predicting the inputs are unsimilar. Gradient-based methods such as Grad-ECLIP fail to recover nonlinear rankings of important tokens, whereas FIxLIP faithfully recovers the optimal subset explanation. exCLIP does not appear in the comparison as its implementation returns an OOM error when querying the larger model ($14\times 14$ image tokens) with texts longer than 24 tokens, which are common in MS COCO.
@@
Table 4: Pointing game recognition for CLIP (ViT-B/16) on ImageNet-1k. Extended Table 1. PGR score for FIxLIP against alternative explanation methods, where a random baseline scores $0.25$. First-order methods, such as Grad-ECLIP and Shapley values, fail to distinguish between multiple objects simultaneously, while second-order methods faithfully recover the appropriate explanation (up to the irreducible non-optimality of the pointing game).
@@
GAME [chefer2021generic] $60_{\pm.11} .$$41_{\pm.04} .$$33_{\pm.02} .$$27_{\pm.01}$
Grad-ECLIP [zhao2024gradeclip] $72_{\pm.13} .$$44_{\pm.04} .$$33_{\pm.01} .$$26_{\pm.01}$
@@ E.1 Guidance on interpreting visualizations of interaction explanations
Finding interesting subsets. FIxLIP explanations allow searching for interesting subsets $M$ like $M_{\mathbf{e},\max,k}$ or $M_{\mathbf{e},\min,k}$ for varying sizes of $k$. We can search the space of interactions and visualize only the explanations for tokens in these subsets, or overlay the masks on top of the input. In Figure 13D, we illustrate $M_{\mathbf{e},\max,7}$ containing tokens related to elephant. Starting the search from the bird token retrieves the $k=4$ subset with information about the bird. While the elephant subset increases the similarity, the bird subset could slightly decrease the similarity, presumably because the model would find the isolated image–text similarity of the elephant clearer.
@@ E.2 Additional examples
Figure 14 demonstrates the utility of conditioning on tokens. Figure 15 shows intriguing examples of interactions in inputs with text written in an image, which is related to the typographic capabilities of large vision–language models [tong2024eyes, gong2025figstep, cao2025scenetap].
@@ E.3 Comparison between CLIP (ViT-B/32) and SigLIP-2 (ViT-B/32)
See Figure 16. Note that these models have different text tokenizers and input image resolutions, i.e. $224\times 224$ vs. $256\times 256$, which is why the images differ slightly.
@@ E.4 Comparison between CLIP (ViT-B/32) and CLIP (ViT-B/16)
Visual examples between CLIP (ViT-B/32) and CLIP (ViT-B/16) are in Figure 17. Notably, in the example including the clock (Fig. 17, row 3), conditioning on the vision token containing the letters T and part of O reveals a strong interaction with another vision token of this word in the “sign” and the text token town. This may suggest the presence of a third-order interaction between these tokens.